OpenStudy (anonymous):

A bowl contains candies of the same size in three flavors: orange, strawberry, and pineapple. If the probability of randomly pulling out an orange candy is 1 over 9, and the probability of randomly pulling out a strawberry candy is 3 over 9, what is the probability of randomly pulling out a pineapple candy? 2/9 3/9 4/9 5/9

Mehek (mehek14):

\(\dfrac{1}{9}+\dfrac{3}{9}\)

Mehek (mehek14):

\(=\dfrac{4}{9}\\ \dfrac{9}{9}-\dfrac{4}{9}=answer\)

Mehek (mehek14):

@amandacross ?

OpenStudy (anonymous):

sorry thank you so much! @Mehek14

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (anonymous):

so the answer would be 5/9 correct? @Mehek14

Mehek (mehek14):

yes

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

can you help me with another problem that i am stuck on? @Mehek14

Mehek (mehek14):

sure ut make another post

Mehek (mehek14):

but*

OpenStudy (anonymous):

Jenny has some tiles in a bag. The tiles are of three different colors: purple, pink, and orange. Jenny randomly pulls a tile out of the bag, records the color, and replaces the tile in the bag. She does this 50 times. The results are recorded in the given table: color of tiles: purple,pink, orange. number of times tile is drawn: 4, 24, 22. What is the experimental probability that Jenny will pull out a pink tile?

Mehek (mehek14):

do you have answer choices?

OpenStudy (anonymous):

22/50 22/28 24/50 4/50

OpenStudy (anonymous):

is it C??

Mehek (mehek14):

from your table it says that pink was drawn 24 times from 50 so \(\dfrac{24}{50}\) experimental probability means the probability of something occurring from what was conducted so all you needed to do was look at the table

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

so its C?

Mehek (mehek14):

yea

OpenStudy (anonymous):

thanks again!

Mehek (mehek14):

yw